Solve the following initial value problem
d^2y/dx^2+dx/dy-2y=-6sin 2x-18cos2x
y(0)=2,y'(0)=2
1
Expert's answer
2020-03-26T16:16:55-0400

"y''+y'-2y=-6 \sin 2x -18 \cos 2x"

Above equation is second order linear nonhomogeneous differential equation with constant coefficients. The general solution of this nonhomogeneous equation is the sum of the general solution "y_0 (x)" of the related homogeneous equation ("y''+y'-2y=0") and a particular solution "y_1(x)" of the nonhomogeneous equation:

"y(x)=y_0(x)+y_1(x)".

The characteristic equation is:

"k^2+k-2=0"

Roots of this equation is "k_1=1" and "k_2=-2", so general solution of the related homogeneous equation is "y_0 (x)=C_1e^x+C_2e^{-2x}".

Let's use method of undetermined coefficient for obtaining of the particular solution of the nonhomogeneous equation "y_1(x)=A \sin 2x +B \cos 2x".

Differentials:

"y_1'(x)=2A \cos 2x -2B \sin 2x"

"y_1''(x)=-4A \sin 2x -4B \cos 2x"

Substitute obtained functions into the initial equation:

"-4A \sin 2x -4B \cos 2x+2A \cos 2x -2B \sin 2x-2(A \sin 2x +B \cos 2x)=-6 \sin 2x -18 \cos 2x"

"-4A \sin 2x -4B \cos 2x+2A \cos 2x -2B \sin 2x-2A \sin 2x -2B \cos 2x=-6 \sin 2x -18 \cos 2x"

"\sin 2x (-4A-2B-2A)+ \cos 2x (-4B +2A -2B)=-6 \sin 2x -18 \cos 2x"

"\sin 2x (-6A-2B)+ \cos 2x (2A-6B)=-6 \sin 2x -18 \cos 2x"

"\begin{cases}
-6A-2B=-6 \\
2A-6B=-18
\end{cases}"

"\begin{cases}
-6A-2B=-6 \\
6A-18B=-54
\end{cases}"

"-20B=-60"

"B=3"

"2A-6 \cdot 3=-18"

"2A-18=-18"

"A=0"

So, "y_1(x)=3 \cos 2x" and "y(x)=C_1e^x+C_2e^{-2x}+3 \cos 2x".

According to the initial value problem:

"y'(x)=C_1e^x-2C_2e^{-2x}-6 \sin 2x"

"\begin{cases}
C_1e^0+C_2e^0+3 \cos 0=2 \\
C_1e^0-2C_2e^0-6 \sin 0=2
\end{cases}"

"\begin{cases}
C_1+C_2+3 =2 \\
C_1-2C_2=2
\end{cases}"

"\begin{cases}
C_1+C_2 =-1 \\
C_1-2C_2=2
\end{cases}"

"3C_2=-3"

"C_2=-1"

"C_1-1 =-1"

"C_1=0"


Finally, "y(x)=-e^{-2x}+3 \cos 2x"